To blow off some steam, you go to lunch and order the fish. Trimethylamine (pKb=4.2) is responsible for the fishy aroma. What would be the pH of a 0.0123 M solution of trimethylamine?
Answer #1

Reaction of trimethyl amine with water is , (CH3)3N + H2O (CH3)3NH + + OH -

Kb for above reaction is Kb = [ (CH3)3NH + ] [ OH - ] / [ (CH3)3N ] = 10 - 4.2  = 6.31 10 -05

Let's use ICE table.

Concentration ( M ) (CH3)3N (CH3)3NH + + OH -
I 0.0123
C - X + X + X
E 0.0123 - X X X

K b = ( X )( X ) / 0.0123  - X = 6.31 10 -05

X 2 / 0.0123  - X = 6.31 10 -05

Trimethyl amine is a weak base. So X is very small as compared to 0.0123. Hence we can write 0.0123 - X0.0123.

X 2 /0.0123  = 6.31 10 -05

X 2 = 6.31 10 -05   0.0123

X 2 = 7.76 10 -07

X = 8.81 10 -04   M

X = 8.81 10 -04 M = [ (CH3)3NH + ] = [ OH - ]

We have, pOH = -log [ OH - ] = -log 8.81 10 -04 = 3.06

Therefore, pH = 14- pOH = 14- 3.06 = 10.94